OpenStudy (anonymous):

One quick algebra question please??? medal rewarded(: The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

What is the formula for determining the area of a rectangle?

OpenStudy (anonymous):

You there?

OpenStudy (midhun.madhu1987):

Area of Rectangle = l X b 280 = l X b ----- (1) Square: Length = (l-5) Breadth = (b+1) But Length = Breadth in case of Square i.e. (l-5) = (b+1) (l-b) = 6 -----(2) From (1), b = 280/l ------(3) Substitute (3) in (2), we get l – 280/l = 6 l^2 – 6l -280 = 0 Solving this Quadratic Equation by suitable methods(Splitting the middle terms), We get l = 20m and b = 14m. LET ME KNOW IF YOU NEED ANY CLARIFICATIONS REGARDING THE SOLVING STEPS... THANKS, Midhun